find the equation of the diameter of a circle whose endpoints are at (6,11) and (-3,1)

Answer by stanbon(75887)

find the equation of the diameter of a circle whose endpoints are at (6,11) and (-3,1)
----------
slope = (11-1)/(6--3) = 10/9
----
Form:: y = mx + b
Solve for "b"::
11 = (10/9)*6 + b
-----
11 = (20/3) + b
b = 13/3
-------------
Equation:
y = (10/9)x + (13/3)
===========
Cheers,
Stan H.
---------------